# Dimensional Uplift in Conformal Field Theories

### v1

@@ Abstract
The n-point functions of any Conformal Field Theory (CFT) in $d$ dimensions can always be interpreted as spatial restrictions of corresponding functions in a higher-dimensional CFT with dimension $d^{\prime}>d$. In particular, when a four-point function in $d$ dimensions has a known conformal block expansion, this expansion can be easily extended to $d^{\prime}=d+2$ due to a remarkable identity among conformal blocks, discovered by Kaviraj, Rychkov, and Trevisani (KRT) as a consequence of Parisi-Sourlas supersymmetry and confirmed to hold in any CFT with $d>1$.
@@ 1 Introduction
Dimensional reduction is a property emerging in some disordered systems. In the random-field Ising model it was observed that the correlation functions reduced to those of the pure Ising model, without the random source, in two fewer dimensions [1, 2]. This was understood by Parisi and Sourlas [3] to be due to a hidden supersymmetric formulation. This dimensional reduction holds near the upper critical dimension of six, but fails in sufficiently low dimensions [4, 5, 6, 7, 8, 9, 10, 11, 12]. This failure is now believed to result from operators that are irrelevant near six dimensions becoming relevant around five dimensions [6, 10, 13].
@@
In the present note we explore such a connection across the space dimensions from the point of view of conformal field theory (CFT). As a result, we give a simple algebraic proof of a surprising exact relation discovered by Kaviraj, Rychkov and Trevisani (KRT) as a direct consequence of the Parisi-Sourlas (PS) supersymmetry and confirmed to hold in any CFT with $d>1$ [10]; this relation expresses any $(d-2)$-dimensional conformal block as a linear combination of five conformal blocks in dimension $d$. We in a sense reverse such a relation by constructing, only utilizing some commutation properties of the Casimir operators, five differential operators, $\Lambda_{i}$, transforming an arbitrary $(d-2)$-dimensional conformal block $g^{(d-2)}_{\Delta,\ell}$ ( describing the contribution to a four-point function of a primary of scaling dimension $\Delta$ and spin $\ell$) into five different conformal blocks in dimenion $d$,
@@
The KRT identity allows one to reconstruct the spectrum of primary operators contributing to a four-point function of a CFT in dimension $d$ once this spectrum is known in a $(d-2)$-dimensional CFT with the same external scalars [23].  [^1] In [23] this result is obtained by assuming the PS supersymmetry.
@@ 2 Notation
In a generic CFT in $d$ dimensions the 4pt function of arbitrary scalars $\mathcal{O}_{i}$ of scaling dimension $\Delta_{i}$ can be parametrized as [24, 25]
@@
The freedom in choosing $d$ has intriguing physical consequences. For instance, if $g(u,v)$ corresponds to the four-point function of a random branched polymer system in dimension $d$, then, due to dimensional reduction, its expansion in $(d-2)$-dimensional conformal blocks provides insight into the spectrum of primary operators associated with the Yang-Lee edge singularity. The dimensional shift by 2 in this case is well understood in terms of Parisi-Sourlas supersymmetry.
@@
Keeping $\delta$ fixed, if the equation $2\delta+2m+2=d$ has a solution, then clearly there exist infinitely many solutions for dimensions $d^{\prime}=d+2k$ with $k=1,2,\dots$
This behaviour seems to be a general feature of local conformal field theories. Actually, as already pointed out in [10], the KRT identity implies that if a scalar primary of scaling dimension $\delta$ has a non-vanishing coupling to the energy-momentum tensor in dimension $d$, then it describes a local CFT in any space dimension $d^{\prime}=d+2k\,\,$ , $(k\in{\mathbb{N}})$ , and this property holds for any real $d>1$.
@@ 3 The proof
We aim to demonstrate a special relationship between the quadratic Casimir oprators in $d$ and $d-2$ dimensions. Specifically, we iteratively construct five differential operators $Y_{i}$ $(i=1,\dots 5)$ where $Y_{1}=1$, $Y_{2}$ is first-order in $z$ and ${\bar{z}}$, $Y_{3}$ and $Y_{4}$ are second-order, and $Y_{5}$ is third-order, such that
@@
where the $\alpha^{j}_{i}$’s are simple linear functions of the Casimir in dimension $d-2$. If we attempt to replace in the rhs $C_{2}[d-2]$ with $C_{2}[d^{\prime}]$ with $d^{\prime}\not=d-2$, the linearity in the $Y_{i}$ is immediately lost, and the algebraic structure of the equation no longer closes.
Basic linear algebra manipulations of this equation allow us to define five operators $\Lambda_{i}$, which promote any conformal block in $d-2$ dimensions to a suitable conformal block in $d$ dimensions - leading to the KRT identity.
@@
$C_{2}[d]\,Y_{1}=Y_{1}\,C_{2}[d-2]+Y_{2}~{},$ (12)
where $Y_{2}=\frac{2z\,{\bar{z}}}{z-{\bar{z}}}\left((1-z)\,\partial_{z}-(1-{\bar{z}})\,\partial_{\bar{z}}\right)$.   We proceed with
@@
this remarkable relation, first discovered by Dolan and Osborn [24] and recently reobtained by Trevisani [23] as a consequence of the PS supersymmetry, tells us that $Y_{3}\,g^{(d-2)}_{\Delta,\ell}$ is proportional to $g^{(d)}_{\Delta+1,\ell-1}$.
Next we obtain
@@
Finally we find
@@
$+Y_{3}\,(2ab-2(d-3)(a+b)+(d-2)^{2}-C_{2}[d-2])~{},$
where $C_{4}$ is the quartic Casimir operator defined by Dolan and Osborn in [24] :
@@
Clearly, when we apply the $C_{2}[d]\,Y_{i}$’s to an arbitrary conformal block in $d-2$ dimensions, these differential operators span a five-dimensional space where the action of $C_{2}[d]$ is represented by a $5\times 5$ matrix:
@@
where ${\sf Y}$ is the column-vector formed by the $Y_{i}$’s, $\kappa$ is the coefficient of $Y_{3}$ in Eq.(16) and $c_{2},c_{4}$ are the eigenvalues of the quadratic and quartic Casimir operators in $d-2$ dimensions:
@@
One can check at once that the five eigenvectors $V_{i},(i=1,2,..,5)$ of the transposed of such a matrix define five differential operators $\Lambda_{i}=V_{i}\cdot Y$ which transform $g^{(d-2)}_{\Delta,\ell}$ into a suitable conformal block in $d$ dimensions.
$\Lambda_{i}\,g^{(d-2)}_{\Delta,\ell}\equiv\Lambda_{\alpha\,\beta}\,\,g^{(d-2)}_{\Delta,\ell}\propto g^{(d)}_{\Delta+\alpha,\ell-\beta}\,,\,\alpha,\beta\in\{0,1,2\},\,\alpha+\beta\,\,\,{\rm even}$ (21)
@@
$\textstyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\frac{(d-4+\ell)(3-d+\Delta)}{2}\,Y_{2}+\frac{(d-4+\ell)(3-d+\Delta)(d+\ell-\Delta+2)}{2}\big{]}\,,$
$\Lambda_{20}$ $=\textstyle~{}n_{20}\big{[}Y_{5}+\frac{2d-6+\ell-\Delta}{2}\,Y_{4}+(3+a+b+\frac{2a\,b}{\ell+\Delta}-\frac{\Delta+\ell}{2})\,Y_{3}-$
@@
$\textstyle~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}~{}+\frac{\ell(d-3-\Delta)}{2}\,Y_{2}+\frac{\ell(d-3-\Delta)(\Delta+\ell-2)}{2}\big{]}\,,$
@@
The normalization coefficients $n_{\alpha\,\beta}$ can be chosen in such a way that $\sum_{i}\Lambda_{i}=1$, which implies of course five linear equations for the $n_{\alpha\,\beta}$’s. Thier solution yields
$n_{00}$ $=\textstyle~{}\frac{-2}{(d-4+2\ell)(d-2\Delta)(d-3+\ell-\Delta)}\,,$
$n_{20}$ $=\textstyle~{}\frac{-2}{(d-4+2\ell)(d-2-2\Delta)(\ell+\Delta-1)}\,,$
@@
$n_{02}$ $=\textstyle~{}\frac{2}{(d-4+2\ell)(d-2-2\Delta)(d-3+\ell-\Delta)}\,,$
$n_{22}$ $=\textstyle~{}\frac{2}{(d-4+2\ell)(d-2-2\Delta)(\ell+\Delta-1)}\,.$
Taking advantage of this normalization we finally obtain the KRT identyty
@@
Similar, simpler dimensional uplift identities for two-point and three-point functions of CFT with boundary [29, 30] and for two-point functions in real projective space [31] have been found. All of them can be reformulated in terms of differential operators transforming a conformal block in dimension $d-2$ into a conformal block in dimension $d$. For instance, in the case of CFT with boundaries, the boundary conformal blocks $f_{bdy}^{(d)}[\Delta,\xi]$ are eigenfunctions of the following quadratic Casimir
@@
[…]
@@ 4 Conclusions
We provided a straightforward algebraic proof of a remarkable identity discovered some time ago by Kaviraj, Rychkov, and Trevisani (KRT). This identity expresses a conformal block contributing to the four-point function in $(d-2)$ dimensions as a linear combination of five conformal blocks in $d$ dimensions. The key mechanism involved the iterative construction of five differential operators that transform linearly under the action of the quadratic Casimir. This, in turn, enabled us to define five differential operators, denoted by $\Lambda_{i}$, which transform any conformal block in $(d-2)$ dimensions into five conformal blocks in $d$ dimensions. These operators can be normalized such that $\sum_{i}\,\Lambda_{i}=1$, from which the KRT identity follows immediately. Similar, simpler identities have been proposed in recent literature, all of which can be reformulated using the present approach. As an example, we explicitly worked out the case of boundary conformal blocks contributing to the two-point functions of boundary conformal field theories.

### v2

@@ Abstract
The n-point functions of any Conformal Field Theory (CFT) in $d$ dimensions can always be interpreted as spatial restrictions of corresponding functions in a higher-dimensional CFT with dimension $d^{\prime}>d$. In particular, when a four-point function in $d$ dimensions has a known conformal block expansion, this expansion can be easily extended to $d^{\prime}=d+2$ due to a remarkable identity among conformal blocks, discovered by Kaviraj, Rychkov and Trevisani (KRT) as a consequence of Parisi-Sourlas supersymmetry and confirmed to hold in any CFT with $d>1$.
@@ 1 Introduction
Dimensional reduction is a property emerging in some disordered systems. In the random-field Ising model it was observed that the correlation functions reduce to those of the pure Ising model, without the random source, in two fewer dimensions [1, 2]. This was understood by Parisi and Sourlas [3] to be due to a hidden supersymmetric formulation. This dimensional reduction holds near the upper critical dimension of six, but fails in sufficiently low dimensions [4, 5, 6, 7, 8, 9, 10, 11, 12]. This failure is now believed to result from operators that are irrelevant near six dimensions becoming relevant around five dimensions [6, 10, 13].
@@
In the present note we explore such a connection across the space dimensions from the point of view of conformal field theory (CFT). As a result, we give a simple algebraic proof of a surprising exact relation discovered by Kaviraj, Rychkov and Trevisani (KRT) as a direct consequence of the Parisi-Sourlas (PS) supersymmetry and confirmed to hold in any CFT with $d>1$ [10]; this relation expresses any (d2)(d-2)-dimensional conformal block as a linear combination of five conformal blocks in dimension $d$. Being these functions defined as eigenfunctions of the Casimir operators, we study some algebraic properties of these operators instead of working directly on the conformal blocks like in the original proof of Ref. [10]. We in a sense reverse the KRT identity by constructing, only utilizing some commutation properties of the Casimir operators, five differential operators, $\Lambda_{i}$, transforming an arbitrary (d2)(d-2)-dimensional conformal block $g^{(d-2)}_{\Delta,\ell}$ ( describing the contribution to a four-point function of a primary of scaling dimension $\Delta$ and spin $\ell$) into five different conformal blocks in dimension $d$,
@@
The KRT identity allows one to reconstruct the spectrum of primary operators contributing to a four-point function of a CFT in dimension $d$ once this spectrum is known in a (d2)(d-2)-dimensional CFT with the same external scalars [23].  [^1] In [23] this result is obtained by assuming the PS supersymmetry.
@@ 2 Notation
In a generic CFT in $d$ dimensions the 4pt function of arbitrary scalars $\mathcal{O}_{i}$ of scaling dimension $\Delta_{i}$ can be parametrised as [24, 25]
@@
The freedom in choosing $d$ has intriguing physical consequences. For instance, if $g(u,v)$ corresponds to the four-point function of a random branched polymer system in dimension $d$, then, due to dimensional reduction, its expansion in (d2)(d-2)-dimensional conformal blocks provides insight into the spectrum of primary operators associated with the Yang-Lee edge singularity. The dimensional shift by 2 in this case is well understood in terms of Parisi-Sourlas supersymmetry.
@@
Keeping $\delta$ fixed, if the equation $2\delta+2m+2=d$ has a solution, then clearly there exist infinitely many solutions for dimensions $d^{\prime}=d+2k$ with $k=1,2,\dots$ .
This behavior seems to be a general feature of local conformal field theories. Actually, as already pointed out in [10], the KRT identity implies that if a scalar primary of scaling dimension $\delta$ has a non-vanishing coupling to the energy-momentum tensor in dimension $d$, then it describes a local CFT in any space dimension $d^{\prime}=d+2k\,\,$ , $(k\in{\mathbb{N}})$ , and this property holds for any real $d>1$.
@@ 3 The proof
We aim to demonstrate a special relationship between the quadratic Casimir operators in $d$ and d2d-2 dimensions. Specifically, we construct iteratively five differential operators $Y_{i}$ $(i=1,\dots 5)$ where $Y_{1}=1$, $Y_{2}$ is first-order in $z$ and $\bar{z}$}, $Y_{3}$ and $Y_{4}$ are second-order, and $Y_{5}$ is third-order, such that
@@
where the $\alpha^{j}_{i}$’s are simple linear functions of the Casimir operators in dimension d2d-2. If we attempt to replace on the right-hand side (RHS) $C_{2}[d-2]$ with $C_{2}[d^{\prime}]$ for $d^{\prime}\neq d-2$, the linearity in the $Y_{i}$ is immediately lost, and additional higher-order differential operators are introduced.
When both sides of Equation (11) act on the conformal block $g^{(d-2)}_{\Delta,\ell}$, the Casimir operators on the RHS can be replaced by their eigenvalues. In this case, the $\alpha^{j}_{i}$’s become real or complex numbers, and Eq. (11) reduces to
$C_{2}[d]\,Y_{i}\,g^{(d-2)}_{\Delta,\ell}=a^{j}_{i}\,Y_{j}\,g^{(d-2)}_{\Delta,\ell}\,\,\,$. (12)
This means that $C_{2}[d]$ acts on the linear space spanned by the five functions $Y_{i}\,g^{(d-2)}_{\Delta,\ell}$ as the matrix $a^{j}_{i}$. In particular, the five left-eigenvectors $V_{i}$ of $a^{j}_{i}$ allow us to define five differential operators
$\Lambda_{i}\equiv\,V_{i}^{j}\,Y_{j}\,\,.$ (13)
These operators promote the conformal block $g^{(d-2)}_{\Delta,\ell}$ to appropriate conformal blocks in $d$ dimensions, as outlined in the Introduction.
@@
$C_{2}[d]\,Y_{1}=Y_{1}\,C_{2}[d-2]+Y_{2}~.$ (14)
Here, $Y_{2}=\frac{2z\,{\bar{z}}}{z-{\bar{z}}}\left((1-z)\,\partial_{z}-(1-{\bar{z}})\,\partial_{\bar{z}}\right)$ is a first-order differential operator.
The action of $C_{2}[d]$ on $Y_{2}$ yields a second-order operator, which we write for convenience as $Y_{3}-Y_{4}$. Specifically, we obtain
@@
this remarkable identity, first discovered by Dolan and Osborn [24] and recently derived by Trevisani [23] as a consequence of the PS supersymmetry, shows that $Y_{3}\,g^{(d-2)}_{\Delta,\ell}$ is proportional to $g^{(d)}_{\Delta+1,\ell-1}$, hence it already corresponds to one of the desired $\Lambda_{i}$.
$C_{2}[d]\,Y_{4}$ yields a third-order operator, which we denote by $Y_{5}$. We find
@@
Finally, applying $C_{2}[d]$ to $Y_{5}$ generates a fourth-order operator. However this operator turns out to be the quartic Casimir operator in (d2)(d-2) dimensions. Specifically we find
@@
$+Y_{3}\,(2ab-2(d-3)(a+b)+(d-2)^{2}-C_{2}[d-2])~.$
Here, $C_{4}$ is the quartic Casimir operator as defined by Dolan and Osborn in [24] :
@@
When Eq.(18) is applied to $g^{d-2}_{\Delta,\ell}$, $C_{4}[d-2]$ is replaced by its eigenvalue, therefore our iterative construction of $Y_{i}$’s terminates at $i=5$ and no higher-order differential operators arise.
The five Eq.s (14-18), when acting on $g^{(d-2)}_{\Delta,\ell}$, can be recast into the matrix form given in (12), namely
@@
where $\sf Y$} is the column-vector formed by the $Y_{i}$’s, $\kappa$ is the coefficient of $Y_{3}$ in Eq.(18) and $c_{2},c_{4}$ are the eigenvalues of the quadratic and quartic Casimir operators in d2d-2 dimensions [24]:
@@
As a consistency check, one can verify that the five eigenvalues of the matrix (20) correspond to the five primary operators in $d$ dimensions contributing to the KRT identity, namely
$\Delta_{i}=\Delta+\alpha\,\,\,\,,\,\,\ell_{i}=\ell-\beta\,\,\,,\alpha,\beta\,\in\{0,1,2\},\,\,\alpha+\beta\,\,\,{\rm even}\,.$ (23)
Following Eq. (13) we can compute the five left-eigenvectors of the matrix and define five differential operators $\Lambda_{i}=V_{i}\cdot Y$ which transform $g^{(d-2)}_{\Delta,\ell}$ into the corresponding conformal block in $d$ dimensions. It is convenient to replace the index $i$ with the two indices $\alpha,\beta$ defined in Eq.(23) which describe the action of $\Lambda_{i}$ on $\Delta$ and $\ell$:
$\Lambda_{i}\,g^{(d-2)}_{\Delta,\ell}\equiv\Lambda_{\alpha\,\beta}\,\,g^{(d-2)}_{\Delta,\ell}\propto g^{(d)}_{\Delta+\alpha,\ell-\beta}\,,\,\alpha,\beta\in\{0,1,2\},\,\alpha+\beta\,\,\,{\rm even}\,.$ (24)
@@
$\textstyle~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+\frac{(d-4+\ell)(3-d+\Delta)}{2}\,Y_{2}+\frac{(d-4+\ell)(3-d+\Delta)(d+\ell-\Delta-2)}{2}\big]\,$,
$\Lambda_{20}$ $=\textstyle~n_{20}\big[Y_{5}+\frac{2d-6+\ell+\Delta}{2}\,Y_{4}+(3+a+b+\frac{2a\,b}{\ell+\Delta}-\frac{\Delta+\ell}{2})\,Y_{3}$-
@@
$\textstyle~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~+\frac{\ell(d-3-\Delta)}{2}\,Y_{2}-\frac{\ell(d-3-\Delta)(\Delta+\ell-2)}{2}\big]\,$,
@@
The normalization coefficients $n_{\alpha\,\beta}$ can be chosen in such a way that $\sum_{i}\Lambda_{i}=1$, which implies of course five linear equations for the $n_{\alpha\,\beta}$’s. Their solution yields
$n_{00}$ $=\textstyle~\frac{-2}{(d-4+2\ell)(d-2-2\Delta)(d-3+\ell-\Delta)}\,$,
$n_{20}$ $=\textstyle~\frac{2}{(d-4+2\ell)(d-2-2\Delta)(\ell+\Delta-1)}\,$,
@@
$n_{02}$ $=\textstyle~\frac{-2}{(d-4+2\ell)(d-2-2\Delta)(d+\ell-1)}\,$,
$n_{22}$ $=\textstyle~\frac{2}{(d-4+2\ell)(d-2-2\Delta)(d-3+\ell-\Delta)}\,.$
Taking advantage of this normalization we finally obtain the KRT identity
@@
It is worth noting that the KRT identity simplify considerably for scalar conformal blocks (i.e. $\ell=0$), where only two terms remain. This simplification can be traced back to Eq.(16), which shows that the operator $Y_{3}$ reduces the spin $\ell$ by one. Since there are no conformal blocks with negative spin, $Y_{3}$ annihilates any scalar conformal block. Furthermore, $Y_{4}$ and $Y_{5}$ are proportional to $Y_{3}$ or its derivatives, so they also annihilate this kind of conformal block. Consequently, in the scalar case, the iterative construction of the $Y_{i}$’s terminates at $i=2$. The two remaining $\Lambda$ operators become in this case
$\Lambda_{00}^{(\ell=0)}=\frac{Y_{2}+d-2-\Delta}{d-2-2\Delta}\,,\,\Lambda_{20}^{(\ell=0)}=\frac{-Y_{2}-\Delta}{d-2-2\Delta}\,,\Lambda^{(\ell=0)}_{00}+\Lambda^{(\ell=0)}_{20}=1\,.$ (29)
Similar, simpler dimensional uplift identities for two-point and three-point functions of CFT with boundary [29, 30] and for two-point functions in real projective space [31] have been found. All of them can be reformulated in terms of differential operators transforming a conformal block in dimension d2d-2 into a conformal block in dimension $d$. For instance, in the case of CFT with boundaries, the boundary conformal blocks $f_{bdy}^{(d)}[\Delta,\xi]$ are eigenfunctions of the following quadratic Casimir operator
@@
where $k$ is a normalization-dependent coefficient. If for $\xi\to\infty$ $f_{bdy}^{(d)}[\Delta,\xi]\simeq\,1/\xi^{\Delta}$, we have $k=\frac{\Delta(\Delta+1)}{4(d-2\Delta-5)(d-2\Delta-3)}$.
@@ 4 Conclusions
We provided a straightforward algebraic proof of a remarkable identity discovered some time ago by Kaviraj, Rychkov and Trevisani (KRT). This identity expresses a conformal block contributing to the four-point function in (d2)(d-2) dimensions as a linear combination of five conformal blocks in $d$ dimensions. The key mechanism involved the iterative construction of five differential operators that transform linearly under the action of the quadratic Casimir. This, in turn, enabled us to define five differential operators, denoted by $\Lambda_{i}$, which transform any conformal block in (d2)(d-2) dimensions into five conformal blocks in $d$ dimensions. These operators can be normalized such that $\sum_{i}\,\Lambda_{i}=1$, from which the KRT identity follows immediately. Similar, simpler identities have been proposed in recent literature, all of which can be reformulated using the present approach. As an example, we explicitly worked out the case of boundary conformal blocks contributing to the two-point functions of boundary conformal field theories.